OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (jdoe0001):

what is the slope of -> y = 2x + 2 ?

OpenStudy (anonymous):

jdoe it should be 2 wouldnt you think. I mean i had trouble with this to but i think the slope of it is 2

OpenStudy (jdoe0001):

yes, it's

OpenStudy (anonymous):

y = 2x + 6 y = −one halfx + 3 y = −one halfx + 6 y = 2x + 3 these are the choices

OpenStudy (jdoe0001):

well, perpendicular lines have NEGATIVE RECIPROCAL slopes what does that mean? since this one has 2 the RECIPROCAL of that is 1/2 the NEGATIVE of that is -1/2

OpenStudy (anonymous):

ok so what would the answer be?

OpenStudy (jdoe0001):

so, now you have, the slope of -1/2 of the line that passes through (6, 3) using the point-slope form, of => \(\huge \bf y-y_1=m(x-x_1)\) m = slope (6, 3) = \(\bf \large x_1, y_1\) use that and your values, and solve for "y" to get your equation :)

OpenStudy (anonymous):

ok so would the answer be y = −one halfx + 6

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

thank you

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

I have another question for you if you dont mind

OpenStudy (jdoe0001):

ok